OpenStudy (anonymous):

Decide if these fractions are indeed equivalent by answering yes or no! 2/5=3/9 5/8=10/16 4/6=3/10 15/45=5/15 7/8=9/12 4/11=3/16 14/28=1/2 3/26=7/15

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{c}{d}\iff ad=bc\] so check that

OpenStudy (anonymous):

for example in number 1, \[2\times 9\neq 5\times 3\] so answer is no

OpenStudy (anonymous):

1.) no 2.) yes 3.) no 4.) yes 5.) no 6.) no 7.) yes 8.) no